OpenStudy (anonymous):

PLZ HELP.. if m and k are positive and 10m^2 k^-1 = 100m, what is m^-1 in terms of k? k/10 k/90 1/10k 1/90k

OpenStudy (anonymous):

I know this has already been answered but I still don't get it.

OpenStudy (anonymous):

Anyone?

OpenStudy (anonymous):

it is 1/10k

OpenStudy (anonymous):

Thanks. The book said that too, but I don't understand how they got from Step 3 to 4, if you know what I mean..

OpenStudy (anonymous):

ie: How do I get to this answer?

OpenStudy (anonymous):

10m^2 k^-1 = 100m 10m k^-1 = 100 (m is cancelled from both sides) 10m = 100k (k is multiplied on both sides) m = 10k (10 is divided from both sides) m^-1 = 1/10k (both sides are reciprocated - m^-1 is equivalent to 1/m)

OpenStudy (anonymous):

10m^2k^-1=100m now we cancel 1 m from both the sides so we have , 10mk^-1=100 divide the equation by 10, we get mk^-1 =10 multiply both sides with k m=10k and here we want 1/m so we dived both the sides of the equation with 1 and we have 1/m=1/10k hope this helps !!

OpenStudy (anonymous):

@luxiliu @Living_dreams Thanks a million to you both! Just another question: When you multiplied k on both sides, what happened to the -1 that was above it?

OpenStudy (anonymous):

\[\frac{ 10m ^{2} }{ k } = 100 m\] \[10m ^{2} = 100 mk\] multiply both sides by 1/100m \[\frac{ 10m ^{2} }{ 100m } = k\] \[\frac{ m }{ 10 } = k \] \[m=10k\] take the reciprocal for both sides \[\frac{ 1 }{ m } = \frac{ 1 }{10k }\] \[m ^{-1} = \frac{ 1 }{10k }\]

OpenStudy (anonymous):

you have k i.e k^1 and u mulitpy with k^-1 so the -1 and 1 makes it k^0 and anything raised to 0 is 1.

OpenStudy (anonymous):

Well, anything to the power of -1 is basically it's reciprocal. Therefore, when you multiply k^-1 with k, you get one, and so it is essentially cancelled out and moved to the other side

OpenStudy (anonymous):

Okay, so k^-1 when multiplied against k (its positive reciprocal) it sort of neutralizes the negative power.. is that about right?

OpenStudy (anonymous):

\[m=10 k \] multiply by m^-1 k^-1 k^-1 = 10 m^-1 k^-1 / 10 = m^-1 >> m * m^-1 = m^(1+(-1)) = m^0 = 1

OpenStudy (anonymous):

@Avva Thank you for your help!

OpenStudy (anonymous):

@Dj2k12 URW :)

OpenStudy (anonymous):

mm there will be no power left so it will be a zero. and anything raised to a zero is always a 1

OpenStudy (anonymous):

"Anything raised to a 0 will always be 1" - I'm writing that one down. @Living_dreams, thanks again for your help.

OpenStudy (anonymous):

@Dj2k12 : Your welcome :) good luck !!